5) You are studying a population of ducks. There are 10 ducks in the population and you are studying how much they weigh in pounds. In this population there are 5 ducks that weigh 5 pounds, and 5 ducks that weigh 4 pounds. You take a sample of 4 ducks of the following weights: 4,4,4,5. If you calculate the sample mean weight of these four ducks, what would be your sampling error?
Answer #1

Population mean = (4*5+5*5)/10 = 4.5

Sample mean = (4+4+4+5)/4=4.25

Sampling error= Population mean - sample mean = 4.5-4.25=0.5